5) A) What is the government spending multiplier? Give a numerical example when the marginal propensity to consume is 0.8.

B) What are the reasons given in your text for why unemployment is higher in the Europe than the US?

10)   A) What variables are you assuming are constant in the above IS curve?

B) Relate PI/PC to the slope of the PPC curve and explain how an decrease in it will affect the supply of investment goods using the concept of opportunity cost to a firm.
Answer #1

Question 5 is solved here :

5) A) Government spending multiplier = 1 / MPS = 1 / 1 - MPC

Delta Y / Delta G = 1 / 1 - MPC

It tells us how much unit change in government spending changes national Income or Y . It gives a multiplying effect .

When MPC = 0.8

Multiplier = 1 / 1 - 0.8 = 5

If G changes by $1 then Y will change by $5.

B) European unemployment has been higher due to lack of labor mobility , frequent recessions , political turbulence , frequent changes in trade pattern etc .